a body push a block of 3 kg with 2.5 acceleration then force will be ?


1
Expert's answer
2021-06-28T14:15:01-0400

Force is given by

"F=m\times a\\
m=3kg\\
a=2.5m/s^2\\
F=3\times 2.5\\
F=7.5N"